OpenStudy (anonymous):

Calculate the flux of the vector field F = 3i + 6j through a square of side 4 lying in the plane x + y + z = 15 oriented away from the origin.

OpenStudy (sidsiddhartha):

\[flux=F.A\] so first try to find out the area vector

OpenStudy (sidsiddhartha):

so i can say that the an upward normal vector is\[A_u=\hat{i}+\hat{j}+\hat{k}\]

OpenStudy (sidsiddhartha):

so the unit normal vector will be \[A_{UN}=\frac{ \hat{i}+\hat{j}+\hat{k} }{ \sqrt{1^2+1^2+1^2} }=\frac{ \hat{i}+\hat{j}+\hat{k} }{ \sqrt{3} }\]

OpenStudy (sidsiddhartha):

ok?

OpenStudy (sidsiddhartha):

so the area vector will be\[A_v=A_{UN}*4=\frac{ 4(\hat{i}+\hat{j}+\hat{k}) }{ \sqrt{3} }\]

OpenStudy (sidsiddhartha):

so \[flux=F.A_v\]

OpenStudy (sidsiddhartha):

so do a little dot product

OpenStudy (anonymous):

what would i do with the =15?

OpenStudy (sidsiddhartha):

\[flux=(3 \hat{i}+4 \hat{j})*\frac{ 4(\hat{i}+\hat{j}+\hat{k}) }{ \sqrt{3}}\]

OpenStudy (sidsiddhartha):

15 is just a constant

OpenStudy (anonymous):

so just (12+16+4)/square root(3)

OpenStudy (sidsiddhartha):

why 12+16+14

OpenStudy (anonymous):

so 32/square root(3)

OpenStudy (anonymous):

wait just (12+16)/square root(3)

OpenStudy (sidsiddhartha):

\[Flux=\frac{ 4 }{ \sqrt{3} }*(\hat{i}+\hat{j}+\hat{k})(3 \hat{i}+4\hat{j})=\frac{ 4 }{ \sqrt{3} }*(3+4)=\frac{ 28 }{ \sqrt{3} }\]

OpenStudy (sidsiddhartha):

yes that should be it !!

OpenStudy (sidsiddhartha):

matches your answer?

OpenStudy (anonymous):

got it thanks

OpenStudy (anonymous):

:)

OpenStudy (sidsiddhartha):

yw!!

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

we did it wrong

OpenStudy (anonymous):

it should be (3i+6j) * 4/square root(3) (i+j+k) right?

OpenStudy (anonymous):

@sidsiddhartha

OpenStudy (sidsiddhartha):

oohh my bad sorry

OpenStudy (sidsiddhartha):

then\[36/\sqrt{3}\] right?

OpenStudy (anonymous):

yeah but its making it wrong :(

OpenStudy (anonymous):

marking it wrong

OpenStudy (anonymous):

what are we doing wrong?

OpenStudy (sidsiddhartha):

i dont know @perl can u check this out?

OpenStudy (sidsiddhartha):

ok u can put a minus ahead of the answer and check it again but not sure it will work or not

OpenStudy (anonymous):

no it didnt work

OpenStudy (sidsiddhartha):

as the vector is oriented away from the origin

OpenStudy (anonymous):

it also marking wrong -36/square root(3)

OpenStudy (sidsiddhartha):

try \[12\sqrt{3}\]

OpenStudy (anonymous):

does "square of side 4 lying in the plane" has something to do with it?

OpenStudy (anonymous):

no the 12 didnt work

OpenStudy (anonymous):

I only have one more try left

OpenStudy (sidsiddhartha):

ok let me tag experts @Kainui

OpenStudy (sidsiddhartha):

@ganeshie8

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

what about (3i+6j)*16(i+j+k)/square root(3)

OpenStudy (anonymous):

because of the square of side 4

OpenStudy (sidsiddhartha):

yes i missed it

OpenStudy (anonymous):

you think so?

OpenStudy (sidsiddhartha):

yes area=16

OpenStudy (anonymous):

so 144/square root(3)

OpenStudy (sidsiddhartha):

yes it should be okay !!

OpenStudy (anonymous):

NO :( I got it wrong

OpenStudy (anonymous):

so what do you think the answer should be?

OpenStudy (anonymous):

can you help me with this one

OpenStudy (anonymous):

@sidsiddhartha

OpenStudy (sidsiddhartha):

what were the options for the last one?

OpenStudy (sidsiddhartha):

@SithsAndGiggles

OpenStudy (anonymous):

for the flux? it didn't have options

OpenStudy (anonymous):

can you help me with the screen shot that i post? please :)

ganeshie8 (ganeshie8):

done with the main question right ?

OpenStudy (sidsiddhartha):

no the answer is'nt correct she says, can u check it @ganeshie8

OpenStudy (anonymous):

it was not the current density. I got it wrong. ;( what should the right answer be? @ganeshie8

OpenStudy (anonymous):

yeah I would like to know what we did wrong

OpenStudy (sidsiddhartha):